# Bridging the Sim-to-Real Gap with Bayesian Inference

### v1

@@ V-C Online RL
Refer to captionpt
Sim-FSVGD
pt
FSVGD
[…]
@@
Refer to captionpt
Sim-FSVGD
pt
FSVGD
ptEpisode 1ptTargetptEpisode 1ptTargetptEpisode 10ptTargetptEpisode 10ptTarget
@@
In Figure 7 we report the learning curves for the racecar in simulation and the real RC car. We observe that in both cases, incorporating the simulation prior leads to considerably faster convergence (ca. 2×2$\times$2 × fewer episodes). Moreover, in Figure 3 we show the realized trajectory for the third episode of the racecar. We see that already after three episodes, Sim-FSVGD learns to drive the car close to the target whereas not incorporating the prior leads to exploration away from the target position. Similar behavior is observed on hardware (c.f., Figure 6), where from episode 1, the policy learned using Sim-FSVGD drives the car close to the target and at episode 10 it is parking the car nearly perfectly. On the contrary, not incorporating the prior leads to no movement of the car at episode 1 and the car driving away from the target at episode 10. This experiment demonstrates that incorporating a simulation prior in model-based RL leads to directed exploration which results in better sample-efficiency.
@@ VII Acknowledgments
This project has received funding from the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545, and the Microsoft Swiss Joint Research Center. Jonas Rothfuss was supported by an Apple Scholars in AI/ML fellowship.

### v2

@@ IV Sim-FSVGD
Estimating Functional Prior from Physical Prior
Input: Measurement distribution $\nu$, Physical prior $\bm{g}$, Parameter distribution $p(\bm{\phi})$.
1:Sample measurement set $\mathbf{X}_{M}\sim\nu$
2:Sample parameters $\bm{\phi}_{j}$ for $j\in\{1,\dots,P\}$ from $p(\bm{\phi})$.
3:Calculate physical model output ${\bm{g}}(\mathbf{X}^{M};\bm{\phi}_{j})$ for $j\in\{1,\dots,P\}$.
4:Calculate empirical mean $\mathbf{\mu}({\mathbf{X}_{M}})$ and variance $\mathbf{\Sigma}({\mathbf{X}_{M}})$ over the $P$ samples.
5:Approximate Gaussian prior $p(\mathbf{h}^{\mathbf{X}_{M}})\sim\mathcal{N}(\mathbf{\mu}({\mathbf{X}_{M}}),\mathbf{\Sigma}({\mathbf{X}_{M}}))$
@@ V-C Online RL
Refer to caption Sim-FSVGD FSVGD $x_{0}$arget
@@
Refer to caption Sim-FSVGD FSVGD Episode 1TargetEpisode 1TargetEpisode 10TargetEpisode 10Target
@@
In Figure 7 we report the learning curves for the racecar in simulation and the real RC car. We observe that in both cases, incorporating the simulation prior leads to considerably faster convergence (ca. 2×2$\times$2 × fewer episodes). Moreover, in Figure 3 we show the realized trajectory for the third episode of the racecar. We see that already after three episodes, Sim-FSVGD learns to drive the car close to the target whereas not incorporating the prior leads to exploration away from the target position. Similar behavior is observed on hardware (c.f., Figure 6), where from episode 1, the policy learned using Sim-FSVGD drives the car close to the target, and at episode 10 it is parking the car nearly perfectly. On the contrary, not incorporating the prior leads to no movement of the car in episode 1 and the car driving away from the target in episode 10. This experiment demonstrates that incorporating a simulation prior in model-based RL leads to directed exploration which results in better sample-efficiency.
@@ VII Acknowledgments
This research was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement no. 815943, the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545, and the Microsoft Swiss Joint Research Center. Jonas Rothfuss was supported by an Apple Scholars in AI/ML fellowship.